OpenStudy (anonymous):

. Given the quadratic function f(x) = x2 – 6x + 9, find a value of x such that f(x) = 25.

OpenStudy (cwtan):

x2-6x+9=25 x2-6x-16=0 solve the quadratic eq

OpenStudy (cwtan):

by using factorize (x-8)(x+2)=0 x-8=0 x=8 x+2=0 x=-2

OpenStudy (anonymous):

im confused

OpenStudy (cwtan):

the question say wan to find the x such that f(x)=25 since f(x)=x2-6x+9 and f(x)=25 we combine it bcome x2-6x+9=25

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

f(x) = x^2 – 6x + 9 = 25 (According to the question) \[x^2 -6x+9-25=0\]\[x^2-6x-16=0\]\[(x-8)(x+2) = 0\](Factorized) x=8 or x=-2

OpenStudy (anonymous):

thanks guys